OpenStudy (anonymous):

Evaluate the Definite Integral from 1 to 3 of (5+1/x+e^x)dx

OpenStudy (jkristia):

What is the antiderivative of 5, 1/x and e^x ? When you have that, call it F(x), then find the value F(3) - F(1)

OpenStudy (anonymous):

antiderivative of 5 would be 5x and for 1/x it would be x^-1 (?) and e^x would just be e^x

OpenStudy (anonymous):

actually 1/x's antiderivative would be ln x

OpenStudy (anonymous):

CAn anyone help?

OpenStudy (e.mccormick):

Yah, \[\int_1^3\left(5+\frac{1}{x}+e^x\right)dx= \left( 5x+\ln x + e^x\right)|_1^3\]

OpenStudy (e.mccormick):

So the next step is to evaluate them on that range. Did you understand what jkristia meant by F(3)-F(1)?

OpenStudy (e.mccormick):

If: \[f(x)=5+\frac{1}{x}+e^x \]then the antiderivative \[F(x)=\left( 5x+\ln x + e^x+C\right)\] When evaluating over a range of: \[\int_a^b\] You solve for \(F(b)-F(a)\), which should have been shown to you with limits, and because no matter what F(x) is, this will invove: polynomial + C - [ polynomial + C] Which can be rewritten as: polynomial - polynomial + C - C And since C-C=0, you can ignore the C. That means in your problem you evaluate: \(F(3)-F(1)\) where you use the C-less form of \(F(x)=\left( 5x+\ln x + e^x\right)\)